C12 c14 98% find the number of. C14 atoms in 12g of sample

C12 c14 98%  find the number of. C14 atoms in 12g of sample

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Radha
 

14 g of carbon -14 atoms contains = 6.023 × 1023 atoms

1g of carbon -14 atom contains =  6.023 × 1023/14

12g of carbon -14 atom contains = (6.023 × 1023 × 12) / 14

 = 5.16 × 1023

 Now its 2% is (5.16 × 1023  × 2 )/100

 =  10.32 × 1021

So, the number of C-14 atoms in 12 g of carbon is 10.32 × 1021

 

 

Regards

Arun (askIITians forum expert)